OpenStudy (anonymous):

Staples sells CD's ($2.00) and boxes of computer paper ($3). If total sales were $1,560 and customers bought 5 times as many CD's as boxes of computer paper, how many CD's were sold

OpenStudy (anonymous):

600 CDs were sold and 120 boxes.

OpenStudy (anonymous):

System of equations here. Therefore you need two equations: 1560 = 2c + 3b c = 5b

OpenStudy (anonymous):

i get it now

OpenStudy (anonymous):

Plug in 5b for c in the first equation: 1560 = 2(5b) + 3b Simplify 1560 = 10b + 3b 1560 = 13b Solve for b by dividing both sides by 13. (b)oxes of paper = 120 c = 5(b) = 5(120) = 600

OpenStudy (anonymous):

That's a little more thorough.

OpenStudy (anonymous):

it helps and is gretaly appreciated

OpenStudy (anonymous):

Not a problem. Take care